OpenStudy (anonymous):

An arc of length 10 meters is formed by a central angle A on a circle of radius 4. The measure of A in degrees (to two decimal places) is _______.

ganeshie8 (ganeshie8):

\(2 \pi r\) is the perimeter which corresponds to a total 360 degrees of central angle

ganeshie8 (ganeshie8):

setup a proportion :- \(\frac{10}{2 \pi r} = \frac{A}{360}\) \(\frac{10}{2 \pi * 4} = \frac{A}{360}\) solve A

OpenStudy (anonymous):

Nice presentation of the problem @ganeshie8

ganeshie8 (ganeshie8):

ty :)

OpenStudy (anonymous):

so is it going to be 143.24?

ganeshie8 (ganeshie8):

Yes, good work !

OpenStudy (anonymous):

awesome thank you

ganeshie8 (ganeshie8):

yw :)